# Certain finite groups whose commuting conjugacy class graph satisfy Hansen-Vukičević conjecture

### v1

@@ 2.1. Zagreb indices of CCC-graphs of certain well-known finite groups
We have $\mathcal{CCC}(G)=K_{2m-2}\sqcup 2K_{2}$. As such, $|v(\mathcal{CCC}(G))|=2m-2+4=2m+2$ and $|e(\mathcal{CCC}(G))|=\binom{2m-2}{2}+2=\frac{(2m-2)(2m-3)}{2}+2=(m-1)(2m-3)+2$. By Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))=(2m-2)(2m-2-1)^{2}+2\times 2(2-1)^{2}=(2m-2)(2m-3)^{2}+4$
and
$M_{2}(\mathcal{CCC}(G))=\frac{(2m-2)(2m-2-1)^{3}}{2}+2\times\frac{2(2-1)^{3}}{2}=(m-1)(2m-3)^{3}+2.$
Therefore
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$ $=\frac{(m-1)(2m-3)^{3}+2}{(m-1)(2m-3)+2}-\frac{(2m-2)(2m-3)^{2}+4}{2m+2}$
$=\frac{(m-1)(2m-3)^{3}+2}{(m-1)(2m-3)+2}-\frac{(m-1)(2m-3)^{2}+2}{m+1}$
$=\frac{8m^{4}-48m^{3}+104m^{2}-96m+32}{(m+1)(m-1)(2m-3)+2(m+1)}$
$=\frac{8m^{3}(m-6)+8m(13m-12)+32}{(m+1)(m-1)(2m-3)+2(m+1)}:=\frac{f(m)}{g(m)}.$
@@
We have $\mathcal{CCC}(G)=2K_{n}\sqcup K_{n(\frac{m}{2}-1)}$. As such, $|v(\mathcal{CCC}(G))|=2n+n(\frac{m}{2}-1)=\frac{1}{2}(mn+2n)$ and $|e(\mathcal{CCC}(G))|=2\binom{n}{2}+\binom{\frac{mn}{2}-n}{2}=\frac{1}{8}(8n(n-1)+(mn-2n)(mn-2n-2))$. By Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))$ $=2n(n-1)^{2}+\left(\frac{mn}{2}-n\right)\left(\frac{mn}{2}-n-1\right)^{2}$
$=\frac{1}{8}\left(16n(n-1)^{2}+(mn-2n)(mn-2n-2)^{2}\right)$
and
$M_{2}(\mathcal{CCC}(G))$ $=2n\frac{(n-1)^{3}}{2}+\left(\frac{mn}{2}-n\right)\frac{\left(\frac{mn}{2}-n-1\right)^{3}}{2}+$
$=\frac{1}{32}\left(32n(n-1)^{3}+(mn-2n)(mn-2n-2)^{3}\right).$
Therefore
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{\frac{1}{32}\left(32n(n-1)^{3}+(mn-2n)(mn-2n-2)^{3}\right)}{\frac{1}{8}(8n(n-1)+(mn-2n)(mn-2n-2))}$
$\hskip 156.49014pt-\frac{\frac{1}{8}\left(16n(n-1)^{2}+(mn-2n)(mn-2n-2)^{2}\right)}{\frac{1}{2}(mn+2n)}$
@@
Note that $g_{1}(m,n)=16\times|e(\mathcal{CCC}(G))|\times|v(\mathcal{CCC}(G))|>0$. We have $f_{1}(4,n)=0$. Also, since $m-2>0$, $m-4>0$ and $mn-4>0$ for all $m\geq 6$ and $n\geq 2$, therefore, $\frac{1}{4}\times\frac{f_{1}(m,n)}{g_{1}(m,n)}\geq 0$. Hence the result follows.
@@
We have $\mathcal{CCC}(G)=K_{2n}\sqcup K_{n(\frac{m}{2}-1)}$. As such, $|v(\mathcal{CCC}(G))|=2n+n(\frac{m}{2}-1)=\frac{1}{2}(mn+2n)$ and $|e(\mathcal{CCC}(G))|=\binom{2n}{2}+\binom{\frac{mn}{2}-n}{2}=\frac{1}{8}(8n(2n-1)+(mn-2n)(mn-2n-2))$. By Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))$ $=2n(2n-1)^{2}+\left(\frac{mn}{2}-n\right)\left(\frac{mn}{2}-n-1\right)^{2}$
$=\frac{1}{8}\left(16n(2n-1)^{2}+(mn-2n)(mn-2n-2)^{2}\right)$
and
$M_{2}(\mathcal{CCC}(G))$ $=2n\frac{(2n-1)^{3}}{2}+\left(\frac{mn}{2}-n\right)\frac{\left(\frac{mn}{2}-n-1\right)^{3}}{2}$
$=\frac{1}{32}\left(32n(2n-1)^{3}+(mn-2n)(mn-2n-2)^{3}\right).$
Therefore
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{\frac{1}{32}\left(32n(2n-1)^{3}+(mn-2n)(mn-2n-2)^{3}\right)}{\frac{1}{8}(8n(2n-1)+(mn-2n)(mn-2n-2))}$
$\hskip 156.49014pt-\frac{\frac{1}{8}\left(16n(2n-1)^{2}+(mn-2n)(mn-2n-2)^{2}\right)}{\frac{1}{2}(mn+2n)}$
@@
Note that $g_{2}(m,n)=16\times|e(\mathcal{CCC}(G))|\times|v(\mathcal{CCC}(G))|>0$. We have $f_{2}(6,n)=0$. Also, since $m-2>0$, $m-6>0$ and $mn+2n-4>0$ for all $m\geq 10$ and $n\geq 2$, therefore, $\frac{1}{4}\times\frac{f_{2}(m,n)}{g_{2}(m,n)}\geq 0$. Hence the result follows.
@@
We have $\mathcal{CCC}(G)=K_{n}\sqcup K_{\frac{n(m-1)}{2}}$. As such, $|v(\mathcal{CCC}(G))|=\frac{mn-n}{2}+n=\frac{1}{2}(mn+n)$ and $|e(\mathcal{CCC}(G))|=\binom{\frac{mn-n}{2}}{2}+\binom{n}{2}=\frac{1}{8}((mn-n)(mn-n-2)+4n(n-1)$. By Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))$ $=\frac{(mn-n)}{2}\left(\frac{mn-n}{2}-1\right)^{2}+n(n-1)^{2}$
$=\frac{1}{8}\left((mn-n)(mn-n-2)^{2}+8n(n-1)^{2}\right)$
and
$M_{2}(\mathcal{CCC}(G))$ $=\frac{(mn-n)}{2}\frac{\left(\frac{mn-n}{2}-1\right)^{3}}{2}+n\frac{(n-1)^{3}}{2}$
$=\frac{1}{32}\left((mn-n)(mn-n-2)^{3}+16n(n-1)^{3}\right).$
Therefore
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{\frac{1}{32}\left((mn-n)(mn-n-2)^{3}+16n(n-1)^{3}\right)}{\frac{1}{8}((mn-n)(mn-n-2)+4n(n-1))}$
$\hskip 156.49014pt-\frac{\frac{1}{8}\left((mn-n)(mn-n-2)^{2}+8n(n-1)^{2}\right)}{\frac{1}{2}(mn+n)}$
@@
Note that $g_{3}(m,n)=16\times|e(\mathcal{CCC}(G))|\times|v(\mathcal{CCC}(G))|>0$. We have $f_{3}(3,n)=0$, $f_{3}(5,n)=192n^{5}-128n^{4}>0$, $f_{3}(7,n)=1536n^{5}-768n^{4}>0$ and $f_{3}(9,n)=5760n^{5}-2304n^{4}>0$ for all $n\geq 1$. Also, since $n(m-6)-4=mn-6n-4>0$, $014m-30>$ 0 and $02m-3>$ 0 for all $m\geq 11$ and $n\geq 1$, therefore, $\frac{1}{4}\times\frac{f_{3}(m,n)}{g_{3}(m,n)}\geq 0$. Hence the result follows. ∎
@@
$=2K_{(p^{m+n-1}-p^{m+n-2})}\sqcup(p^{n}-p^{n-1})K_{(p^{m}-p^{m-1})}$. As such, $|v(\mathcal{CCC}(G))|=2(p^{m+n-1}-p^{m+n-2})+(p^{n}-p^{n-1})(p^{m}-p^{m-1})=p^{m+n}-p^{m+n-2}$ and $|e(\mathcal{CCC}(G))|=2\binom{p^{m+n-1}-p^{m+n-2}}{2}+(p^{n}-p^{n-1})\binom{p^{m}-p^{m-1}}{2}$ $=\frac{1}{2}\left(2p^{2m+2n-2}-4p^{2m+2n-3}+2p^{2m+2n-4}+p^{2m+n}-3p^{2m+n-1}\right.\\ \left.+3p^{2m+n-2}-p^{2m+n-3}-p^{m+n}+p^{m+n-2}\right)$. By Theorem 2.1, we have
@@
We have $\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{f(p,m,n)}{g(p,m,n)}$, where
@@ 2.2. Zagreb indices of CCC-graphs of groups with a given central quotient
In this section, we consider the CCC-graphs of the groups $G$ when $G/Z(G)$ is isomorphic to $D_{2m}$, $\mathbb{Z}_{p}\times\mathbb{Z}_{p}$, Frobenious group of order $pq$ and $p^{2}q$ or any group of order $p^{3}$ for any two primes $p$ and $q$. The structures of CCC-graphs of groups when its central quotient is isomorphic to a dihedral group have been obtained in [25, Theorem 1.2]. Using which we get the following result.
@@
We have $\mathcal{CCC}(G)=K_{\frac{x(m-1)}{2}}\sqcup 2K_{\frac{x}{2}}$. As such, $|v(\mathcal{CCC}(G))|=\frac{x(m-1)}{2}+2\times\frac{x}{2}=\frac{x(m+1)}{2}$ and $|e(\mathcal{CCC}(G))|=\binom{\frac{mx-x}{2}}{2}+2\binom{\frac{x}{2}}{2}=\frac{1}{8}((mx-x)(mx-x-2)+2x(x-2))$. By Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))$ $=\frac{(mx-x)}{2}\left(\frac{mx-x}{2}-1\right)^{2}+2\times\frac{x}{2}\left(\frac{x}{2}-1\right)^{2}$
$=\frac{1}{8}\left((mx-x)(mx-x-2)^{2}+2x(x-2)^{2}\right)$
and
$M_{2}(\mathcal{CCC}(G))$ $=\frac{(mx-x)}{2}\frac{\left(\frac{mx-x}{2}-1\right)^{3}}{2}+2\times\frac{x}{2}\frac{\left(\frac{x}{2}-1\right)^{3}}{2}$
$=\frac{1}{32}\left((mx-x)(mx-x-2)^{3}+2x(x-2)^{3}\right).$
Therefore
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{\frac{1}{32}\left((mx-x)(mx-x-2)^{3}+2x(x-2)^{3}\right)}{\frac{1}{8}((mx-x)(mx-x-2)+2x(x-2))}$
$\hskip 156.49014pt-\frac{\frac{1}{8}\left((mx-x)(mx-x-2)^{2}+2x(x-2)^{2}\right)}{\frac{1}{2}x(m+1)}$
$=\frac{1}{4}\times\frac{(mx-x)(mx-x-2)^{3}+2x(x-2)^{3}}{(mx-x)(mx-x-2)+2n(n-2)}-\frac{(mx-x)(mx-x-2)^{2}+2x(x-2)^{2}}{mx+x}$
$=\frac{1}{4}\times\frac{m^{3}x^{3}-4m^{2}x^{3}+4mx^{3}-4m^{2}x^{2}+16mx^{2}-16x^{2}}{mx(mx+x)(mx-2x-2)}$
$=\frac{1}{4}\times\frac{x^{2}(m-2)^{2}(mx-4)}{mx(mx+x)(mx-2x-2)}>0$
@@
We have $\mathcal{CCC}(G)=K_{\frac{x(m-1)}{2}}\sqcup K_{x}$. As such, $|v(\mathcal{CCC}(G))|=\frac{mx-x}{2}+x=\frac{1}{2}(mx+x)$ and $|e(\mathcal{CCC}(G))|=\binom{\frac{mx-x}{2}}{2}+\binom{x}{2}=\frac{1}{8}((mx-x)(mx-x-2)+4x(x-1))$. By Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))$ $=\frac{(mx-x)}{2}\left(\frac{mx-x}{2}-1\right)^{2}+x(x-1)^{2}$
$=\frac{1}{8}\left((mx-x)(mx-x-2)^{2}+8x(x-1)^{2}\right)$
and
$M_{2}(\mathcal{CCC}(G))$ $=\frac{(mx-x)}{2}\frac{\left(\frac{mx-x}{2}-1\right)^{3}}{2}+x\frac{(x-1)^{3}}{2}$
$=\frac{1}{32}\left((mx-x)(mx-x-2)^{3}+16x(x-1)^{3}\right).$
Therefore
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{\frac{1}{32}\left((mx-x)(mx-x-2)^{3}+16x(x-1)^{3}\right)}{\frac{1}{8}((mx-x)(mx-x-2)+4x(x-1))}$
$\hskip 156.49014pt-\frac{\frac{1}{8}\left((mx-x)(mx-x-2)^{2}+8x(x-1)^{2}\right)}{\frac{1}{2}(mx+x)}$
@@
Note that $g(m,x)=16\times|e(\mathcal{CCC}(G))|\times|v(\mathcal{CCC}(G))|>0$. We have $f(3,x)=0$, $f(5,x)=192x^{5}-128x^{4}>0$, $f(7,x)=1536x^{5}-768x^{4}>0$ and $f(9,x)=5760x^{5}-2304x^{4}>0$ for all $x\geq 1$. Also, since $x(m-6)-4=mx-6x-4>0$, $014m-30>$ 0 and $02m-3>$ 0 for all $m\geq 11$ and $x\geq 1$, we have $\frac{1}{4}\frac{f(m,x)}{g(m,x)}\geq 0$. Hence the result follows. ∎
The structures of CCC-graphs of groups when its central quotient is isomorphic to $\mathbb{Z}_{p}\times\mathbb{Z}_{p}$ (for any prime $p$) have been obtained in [26, Theotem 3.1]. Therefore, we get the following result.
@@
We have $\mathcal{CCC}(G)=(p+1)K_{\frac{x(p-1)}{p}}$. As such, $|v(\mathcal{CCC}(G))|=(p+1)\times\frac{x(p-1)}{p}=\frac{x(p^{2}-1)}{p}$ and $|e(\mathcal{CCC}(G))|=(p+1)\binom{\frac{x(p-1)}{p}}{2}=\frac{x(p^{2}-1)}{2p^{2}}(px-x-p)$. By Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))$ $=(p+1)\frac{x(p-1)}{p}\left(\frac{x(p-1)}{p}-1\right)^{2}$
$=\frac{x(p^{2}-1)}{p^{3}}\left(p^{2}x^{2}-2px^{2}+x^{2}-2p^{2}x+2px+p^{2}\right)$
and
$M_{2}$ $(\mathcal{CCC}(G))=(p+1)\frac{x(p-1)}{p}\frac{\left(\frac{x(p-1)}{p}-1\right)^{3}}{2}$
$=\frac{x(p^{2}-1)}{2p^{4}}\left(p^{3}x^{3}-3p^{2}x^{3}+3px^{3}-x^{3}-3p^{3}x^{2}+6p^{2}x^{2}-3px^{2}+3p^{3}x-3p^{2}x-p^{3}\right).$
Now
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}$ $=\frac{p^{3}x^{3}-3p^{2}x^{3}+3px^{3}-x^{3}-3p^{3}x^{2}+6p^{2}x^{2}-3px^{2}+3p^{3}x-3p^{2}x-p^{3}}{p^{2}(px-x-p)}$
$=\frac{(px-x-p)(p^{2}x^{2}-2px^{2}+x^{2}-2p^{2}x+2px+p^{2})}{p^{2}(px-x-p)}$
$=\frac{p^{2}x^{2}-2px^{2}+x^{2}-2p^{2}x+2px+p^{2}}{p^{2}}=\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}.$

### v2

@@ 2.1. Zagreb indices of CCC-graphs of certain well-known finite groups
We have $\mathcal{CCC}(G)=K_{2m-2}\sqcup 2K_{2}$. Therefore, by Theorem 2.1, we get the expressions for $M_{1}(\mathcal{CCC}(G))$ and $M_{2}(\mathcal{CCC}(G))$.
Note that $|v(\mathcal{CCC}(G))|=2m-2+4=2m+2$ and $|e(\mathcal{CCC}(G))|=\binom{2m-2}{2}+2=\frac{(2m-2)(2m-3)}{2}+2=(m-1)(2m-3)+2$. Therefore,
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}=\frac{(m-1)(2m-3)^{3}+2}{(m-1)(2m-3)+2}\text{ and }\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{(m-1)(2m-3)^{2}+2}{m+1}.$
Hence,
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$ $=\frac{8m^{3}(m-6)+8m(13m-12)+32}{(m+1)(m-1)(2m-3)+2(m+1)}:=\frac{f(m)}{g(m)}.$
@@
We have $\mathcal{CCC}(G)=2K_{n}\sqcup K_{n(\frac{m}{2}-1)}$. Therefore, by Theorem 2.1, we get the expressions for $M_{1}(\mathcal{CCC}(G))$ and $M_{2}(\mathcal{CCC}(G))$.
Note that $|v(\mathcal{CCC}(G))|=2n+n(\frac{m}{2}-1)=\frac{1}{2}(mn+2n)$ and $|e(\mathcal{CCC}(G))|=2\binom{n}{2}+\binom{\frac{mn}{2}-n}{2}=\frac{1}{8}(8n(n-1)+(mn-2n)(mn-2n-2))$. Therefore,
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}=\frac{32n(n-1)^{3}+(mn-2n)(mn-2n-2)^{3}}{4(8n(n-1)+(mn-2n)(mn-2n-2))}\text{ and }$
$\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{16n(n-1)^{2}+(mn-2n)(mn-2n-2)^{2}}{4(mn+2n)}.$
Hence,
@@
Note that $g_{1}(m,n)=16\times|e(\mathcal{CCC}(G))|\times|v(\mathcal{CCC}(G))|>0$. We have $f_{1}(4,n)=0$. Also, since $m-2>0$, $m-4>0$ and $mn-4>0$ for all $m\geq 6$ and $n\geq 2$, we have $\frac{1}{4}\times\frac{f_{1}(m,n)}{g_{1}(m,n)}\geq 0$. Hence the result follows.
@@
We have $\mathcal{CCC}(G)=K_{2n}\sqcup K_{n(\frac{m}{2}-1)}$. Therefore, by Theorem 2.1, we get the expressions for $M_{1}(\mathcal{CCC}(G))$ and $M_{2}(\mathcal{CCC}(G))$.
Note that $|v(\mathcal{CCC}(G))|=2n+n(\frac{m}{2}-1)=\frac{1}{2}(mn+2n)$ and $|e(\mathcal{CCC}(G))|=\binom{2n}{2}+\binom{\frac{mn}{2}-n}{2}=\frac{1}{8}(8n(2n-1)+(mn-2n)(mn-2n-2))$. Therefore,
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}=\frac{32n(2n-1)^{3}+(mn-2n)(mn-2n-2)^{3}}{4(8n(2n-1)+(mn-2n)(mn-2n-2))}\text{ and }$
$\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{16n(2n-1)^{2}+(mn-2n)(mn-2n-2)^{2}}{4(mn+2n)}.$
Hence,
@@
Note that $g_{2}(m,n)=16\times|e(\mathcal{CCC}(G))|\times|v(\mathcal{CCC}(G))|>0$. We have $f_{2}(6,n)=0$. Also, since $m-2>0$, $m-6>0$ and $mn+2n-4>0$ for all $m\geq 10$ and $n\geq 2$, we have $\frac{1}{4}\times\frac{f_{2}(m,n)}{g_{2}(m,n)}\geq 0$. Hence the result follows.
@@
We have $\mathcal{CCC}(G)=K_{n}\sqcup K_{\frac{n(m-1)}{2}}$. Therefore, by Theorem 2.1, we get the expressions for $M_{1}(\mathcal{CCC}(G))$ and $M_{2}(\mathcal{CCC}(G))$.
Note that $|v(\mathcal{CCC}(G))|=\frac{mn-n}{2}+n=\frac{1}{2}(mn+n)$ and $|e(\mathcal{CCC}(G))|=\binom{\frac{mn-n}{2}}{2}+\binom{n}{2}=\frac{1}{8}((mn-n)(mn-n-2)+4n(n-1)$. Therefore,
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}=\frac{(mn-n)(mn-n-2)^{3}+16n(n-1)^{3}}{4((mn-n)(mn-n-2)+4n(n-1))}\text{ and }$
$\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{(mn-n)(mn-n-2)^{2}+8n(n-1)^{2}}{4(mn+n)}.$
Hence,
@@
Note that $g_{3}(m,n)=16\times|e(\mathcal{CCC}(G))|\times|v(\mathcal{CCC}(G))|>0$. We have $f_{3}(3,n)=0$, $f_{3}(5,n)=192n^{5}-128n^{4}>0$, $f_{3}(7,n)=1536n^{5}-768n^{4}>0$ and $f_{3}(9,n)=5760n^{5}-2304n^{4}>0$ for all $n\geq 1$. Also, since $n(m-6)-4=mn-6n-4>0$, $014m-30>$ 0 and $02m-3>$ 0 for all $m\geq 11$ and $n\geq 1$, we have $\frac{1}{4}\times\frac{f_{3}(m,n)}{g_{3}(m,n)}\geq 0$. Hence, the result follows. ∎
@@
$=2K_{(p^{m+n-1}-p^{m+n-2})}\sqcup(p^{n}-p^{n-1})K_{(p^{m}-p^{m-1})}$. Therefore, by Theorem 2.1, we have
@@
Note that $|v(\mathcal{CCC}(G))|=2(p^{m+n-1}-p^{m+n-2})+(p^{n}-p^{n-1})(p^{m}-p^{m-1})=p^{m+n}-p^{m+n-2}$ and $|e(\mathcal{CCC}(G))|=2\binom{p^{m+n-1}-p^{m+n-2}}{2}+(p^{n}-p^{n-1})\binom{p^{m}-p^{m-1}}{2}$ $=\frac{1}{2}\left(2p^{2m+2n-2}-4p^{2m+2n-3}+\right.\\ 2p^{2m+2n-4}+p^{2m+n}-3p^{2m+n-1}\left.+3p^{2m+n-2}-p^{2m+n-3}-p^{m+n}+p^{m+n-2}\right)$. We have
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{f(p,m,n)}{g(p,m,n)}$, where
@@ 2.2. Zagreb indices of CCC-graphs of groups with a given central quotient
In this section, we consider the CCC-graphs of the groups $G$ when $G/Z(G)$ is isomorphic to $D_{2m}$, $\mathbb{Z}_{p}\times\mathbb{Z}_{p}$, Frobenious group of order $pq$ and $p^{2}q$ or any group of order $p^{3}$ for any two primes $p$ and $q$. The structures of CCC-graphs of groups when its central quotient is isomorphic to a dihedral group have been obtained in [28, Theorem 1.2]. Using which we get the following result.
@@
We have $\mathcal{CCC}(G)=K_{\frac{x(m-1)}{2}}\sqcup 2K_{\frac{x}{2}}$. Therefore, by Theorem 2.1, we get the expressions for $M_{1}(\mathcal{CCC}(G))$ and $M_{2}(\mathcal{CCC}(G))$.
Note that $|v(\mathcal{CCC}(G))|=\frac{x(m-1)}{2}+2\times\frac{x}{2}=\frac{x(m+1)}{2}$ and $|e(\mathcal{CCC}(G))|=\binom{\frac{mx-x}{2}}{2}+2\binom{\frac{x}{2}}{2}=\frac{1}{8}((mx-x)(mx-x-2)+2x(x-2))$. Therefore,
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}=\frac{(mx-x)(mx-x-2)^{3}+2x(x-2)^{3}}{4((mx-x)(mx-x-2)+2x(x-2))}\text{ and }$
$\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{(mx-x)(mx-x-2)^{2}+2x(x-2)^{2}}{4x(m+1)}.$
Hence,
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$ $=\frac{1}{4}\times\frac{x^{2}(m-2)^{2}(mx-4)}{mx(mx+x)(mx-2x-2)}>0$
@@
We have $\mathcal{CCC}(G)=K_{\frac{x(m-1)}{2}}\sqcup K_{x}$. Therefore, by Theorem 2.1, we get the expressions for $M_{1}(\mathcal{CCC}(G))$ and $M_{2}(\mathcal{CCC}(G))$.
Note that $|v(\mathcal{CCC}(G))|=\frac{mx-x}{2}+x=\frac{1}{2}(mx+x)$ and $|e(\mathcal{CCC}(G))|=\binom{\frac{mx-x}{2}}{2}+\binom{x}{2}=\frac{1}{8}((mx-x)(mx-x-2)+4x(x-1))$. Therefore,
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}=\frac{(mx-x)(mx-x-2)^{3}+16x(x-1)^{3}}{4((mx-x)(mx-x-2)+4x(x-1))}\text{ and }$
$\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{(mx-x)(mx-x-2)^{2}+8x(x-1)^{2}}{4(mx+x)}.$
Hence,
@@
Note that $g(m,x)=16\times|e(\mathcal{CCC}(G))|\times|v(\mathcal{CCC}(G))|>0$. We have $f(3,x)=0$, $f(5,x)=192x^{5}-128x^{4}>0$, $f(7,x)=1536x^{5}-768x^{4}>0$ and $f(9,x)=5760x^{5}-2304x^{4}>0$ for all $x\geq 1$. Also, since $x(m-6)-4=mx-6x-4>0$, $014m-30>$ 0 and $02m-3>$ 0 for all $m\geq 11$ and $x\geq 1$, we have $\frac{1}{4}\frac{f(m,x)}{g(m,x)}\geq 0$. Hence, the result follows. ∎
The structures of CCC-graphs of groups when its central quotient is isomorphic to $\mathbb{Z}_{p}\times\mathbb{Z}_{p}$ (for any prime $p$) have been obtained in [29, Theotem 3.1]. Therefore, we get the following result.
@@
We have $\mathcal{CCC}(G)=(p+1)K_{\frac{x(p-1)}{p}}$. Therefore, by Theorem 2.1, we get the expressions for $M_{1}(\mathcal{CCC}(G))$ and $M_{2}(\mathcal{CCC}(G))$.
Note that $|v(\mathcal{CCC}(G))|=(p+1)\times\frac{x(p-1)}{p}=\frac{x(p^{2}-1)}{p}$ and $|e(\mathcal{CCC}(G))|=(p+1)\binom{\frac{x(p-1)}{p}}{2}=\frac{x(p^{2}-1)}{2p^{2}}(px-x-p)$. We have
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}$ $=\frac{p^{2}x^{2}-2px^{2}+x^{2}-2p^{2}x+2px+p^{2}}{p^{2}}=\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}.$